# Cross Validation in Stochastic Analytic Continuation

### v1

@@ Abstract
Stochastic Analytic Continuation (SAC) of Quantum Monte Carlo (QMC) imaginary-time correlation function data is a valuable tool in connecting many-body models to experiments. Recent developments of the SAC method have allowed for spectral functions with sharp features, e.g. narrow peaks and divergent edges, to be resolved with unprecedented fidelity. Often times, it is not known what exact sharp features are present a priori, and, due to the ill-posed nature of the analytic continuation problem, multiple spectral representations may be acceptable. In this work, we borrow from the machine learning and statistics literature and implement a cross validation technique to provide an unbiased method to identify the most likely spectrum. We show examples using imaginary-time data generated by QMC simulations and synthetic data generated from artificial spectra. Our procedure, which can be considered a form of “model selection,” can be applied to a variety of numerical analytic continuation methods, beyond just SAC.
@@ I Introduction
Dynamic response functions (spectral functions) provide a key link between the quantum Monte Carlo (QMC) simulations of many-body systems and their experimental counterparts. While it is intractable to measure these spectral function directly, one can instead preform numerical analytic continuation of imaginary-time correlation functions to real frequency. However, this process is what is considered an “ill-posed” problem, meaning that there may be many viable solutions for the spectral function fitting a set of QMC generated data. The Maximum Entropy method (MEM) [1] has been the traditional tool used to tackle this problem and has produced many important results [2]. However, MEM has its limitations, being able only to reproduce broad spectral features, unless prior information about more complicated structure is provided to the algorithm.
@@
An alternative approach to MEM is to instead rely on the stochastic averaging of many solutions for the spectral function that fit the QMC data well. The Stochastic Analytic Continuation (SAC) method [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 12, 14, 15], also referred to as the average spectrum method, achieves this by regularizing the average spectra using a fictitious temperature, $\Theta$, which mediates the competition between the minimization of the goodness-of-fit $\chi^{2}$ with respect to the QMC data and the favoring of spectra with the high configurational entropy. It is now known that this procedure is equivalent to the MEM when the number of sampled degrees of freedom of the spectrum is large and when the two resulting spectra are compared at the same $\chi^{2}$ value (although the definition of the entropy may need to be adjusted depending on what degrees of freedom are sampled) [5, 14, 15]. Recent developments in the SAC method have also allowed for the resolution of spectra with sharp features, such as narrow quasi-particle peaks and edges with power-law singularities, by imposing constraints on the sampling space of the spectral function [10, 14].
@@
In both the plain, “unconstrained” version of SAC, and the newly developed “constrained” SAC sampling schemes, the average $\chi^{2}$ of the sampled spectra is used to optimize various parameters and gauge the ability of the spectral representation to describe the data. While these methods that rely on the average $\chi^{2}$ alone have been shown to produce reliable and consistent results, the goal of this study is to provide further confirmation for these simple schemes by taking advantage of concepts from the machine learning and statistics literature. We do this by implementing a cross validation procedure [16, 17].
@@
Cross validation is a tool used to compare how well different models describe a set of data. This is generally achieved by splitting the data into mutually exclusive sets, a training, or sampling, set and a series of testing, or validation, sets. A model is trained on the sampling set, typically by optimizing some type of loss function. After training, the ability of the optimized model to describe the validation data sets is determined using this same loss function. This can provide information about the suitability of the model to describe the data, as well as the level of over-fitting to noise in the sampling data set.
@@
Cross validation has a natural extension to SAC. The imaginary-time correlation function data, calculated using QMC simulations, is typically stored as individual bins, which can be split up into sampling and validation sets. Here, the goodness-of-fit $\chi^{2}$ is used as the loss function to “train” and to “validate” the resulting spectral function. The concept of using cross validation within SAC was first implemented by Efremkin et al. [18], whose work we draw on and significantly expand upon. Here, we preform an exhaustive study using both synthetic and real QMC data.
@@
In Sec. II we describe the SAC procedure generally, and in Sec. III we explain our crossvalidation scheme in detail. In Sec. IV.1 and  IV.2 we implement cross validation using the unconstrained SAC method and demonstrate how it is used to determine the optimal sampling temperature, which we compare to the simple criterion previously proposed by Ref. [14]. In Sec. V.1 and  V.2, we show how cross validation can be used in practice to determine which SAC sampling parameterization produces the most statistically likely spectrum. In Sec. V.1 we preform a test using QMC generated data for the $S=1/2$ antiferromagnetic (AFM) Heisenberg model, where it is known that the spectral function for the $O=S_{q}^{z}$ operator contains a power-law divergent edge. In Sec. V.2 we preform a preliminary test on the $S=1/2$ AFM Heisenberg model with long-range interactions [19, 20], a model where the exact features of the spectral function are unknown.
@@ II SAC Procedure
We consider the imaginary-time correlation function of a bosonic operator $O$ at inverse temperature $1/T=\beta$, evaluated at time points $\tau\in\left[0,\beta/2\right]$,
@@
Here, the imaginary-time dependence is defined in the Heisenberg picture, taking $\hbar=1$,
@@
The corresponding spectral function can be defined in terms of the eigenstates $|n\rangle$ of the Hamiltonian $H$ describing the system,
@@
For the cross validation procedure, we consider $N_{B}$ bins of imaginary-time correlation function data evaluated at a set of $N_{\tau}$ points $\tau_{i}$, $\{G^{b}(\tau_{i})\}$, where $b=1,2,\ldots,N_{B}$ and $i=1,2,\ldots,N_{\tau}$. The $N_{B}$ bins are then split into $K+1$ mutually exclusive sets of equal length $N_{k}={N_{B}}/\left({K+1}\right)$, $\{G^{b}(\tau_{i})\}_{k}$. For a single cross validation run, one of the $K+1$ groups will act as the sampling data set, while the other $K$ groups will be used for cross validation. For each group, we compute the average
@@
Because the statistical errors of different $\tau$ points are correlated, the full characterization of the error, and thus the calculation of $\chi^{2}$, requires the covariance matrix [1]. We use the bootstrap method to calculate the covariance matrices for each set,
$C_{k}^{ij}=\frac{1}{M}\sum_{m=1}^{M}(G_{k}^{m}(\tau_{i})-\bar{G}_{k}(\tau_{i}))(G_{k}^{m}(\tau_{j})-\bar{G}_{k}(\tau_{j}))$ (7)
where $G_{k}^{m}(\tau_{i})$ is the average of a bootstrap sample of $N_{k}$ randomly chosen bins among the $N_{k}$ bins in the $k$th data set, and $M$ is the total number of bootstrap samples, typically much larger than $N_{B}$.
Given $S(\omega)$, the corresponding $G(\tau_{i})$ values are computed according to Eq. (4) and the $\chi^{2}$ with respect to the QMC data is given by
$\chi^{2}_{k}=\sum_{i,j}^{N_{\tau}}[G(\tau_{i})-\bar{G}_{k}(\tau_{i})]\left[C_{k}^{-1}\right]^{ij}[G(\tau_{j})-\bar{G}_{k}(\tau_{j})].$ (8)
Alternatively, one can compute $\chi^{2}_{k}$ is the eigenbasis of the covariance matrix:
@@
where $\left(\sigma_{k}(\tau_{i})\right)^{2}$ are the eigenvalues of the covariance matrix (and both $G(\tau_{i})$ and $\bar{G}_{k}$ are now computed in this eigenbasis).
In a SAC run, the spectra $S(\omega)$ are importance sampled with a weight
@@
The fictitious temperature $\Theta$ is gradually reduced, as in simulated annealing, until a minimum is reached. In the limit $\Theta\to 0$, the sampled spectrum purely minimizes $\chi^{2}$, which will not reproduce the most probable spectrum when noisy data is used. Instead, it has been proposed to use an elevated temperature, which gives the optimal balance of entropy and goodness-of-fit, corresponding to the value where
$\langle\chi^{2}(\Theta)\rangle=\chi^{2}_{\rm min}+a\sigma_{\chi^{2}}$ (11)
where $\sigma_{\chi^{2}}=\sqrt{2\chi^{2}_{\rm min}}$ is used as a proxy for the standard deviation of the $\chi^{2}$ distribution that the value of $\langle\chi^{2}\rangle$ follows and $a$ is an order one number (typically $a=1/2$ is used). This criterion, motivated by the properties of the $\chi^{2}$ distribution, has been shown to produce reliable results, reproducing spectra generated from synthetic QMC $G(\tau)$ data with high fidelity [14].
@@ III Cross Validation Procedure
which factors into three terms:
@@
$X_{3}=\frac{1}{K}\sum_{k=1}^{K}\sum_{i=1}^{N_{\tau}}\frac{2\left(\bar{G}_{k}(\tau_{i})-G_{\rm ex}(\tau_{i})\right)\left({G}_{\rm ex}(\tau_{i})-{G}(\tau_{i})\right)}{\sigma^{2}_{k},(\tau_{i})}.$ (15c)
Here, $X_{1}$ contains purely information about how well the estimated $G(\tau)$, calculated from the sampled $S(\omega)$, describes the exact correlation function. This will vary as we change $\Theta$, or any other parameter used in the SAC program, and will be minimized by the most statistically likely spectral function, given input data $\bar{G}_{0}(\tau)$ and $C_{0}^{ij}$.
$X_{2}$ is the the $\chi^{2}$ value for the validation data with respect to the underlying data $G_{\rm ex}$, averaged over each of the $K$ validation data sets. The value of $X_{2}$ will follow the $\chi^{2}$ distribution with $N_{\rm dof}=N_{\tau}$, and will have a mean of $E\left[X_{2}\right]=N_{\rm dof}=N_{\tau}$ and a variance of $V\left[X_{2}\right]=2N_{\rm dof}=2N_{\tau}$. This term is independent of any SAC parameter and will simply contribute a background value of $N_{\tau}$, in the limit of large $K$.
$X_{3}$ contains both information about the accuracy of $G(\tau)$ with respect to ${G}_{\rm ex}(\tau)$ and the deviation of the $\bar{G}_{k}(\tau)$ with respect to ${G}_{\rm ex}(\tau)$. However, we expect that fluctuations of $\bar{G}_{k}(\tau)$ will follow a Gaussian distribution with a mean of zero, so in the limit of large $K$, this term should also have a mean of zero. While, $X_{3}$ does depend on the SAC parameters (e.g. $\Theta$), as well as on the covariance of each validation data set, we have found that the dominant contribution is indeed from the Gaussian fluctuations of $\bar{G}_{k}(\tau)$, so this term can be approximated as zero when $K$ is large.
@@
$\chi^{2}_{\mathrm{val}}/N_{\tau}\approx\frac{1}{K}\sum_{k=1}^{K}\sum_{i=1}^{N_{\tau}}\left(\frac{G(\tau_{i})-{G}_{\rm ex}(\tau_{i})}{\sigma_{k}(\tau_{i})}\right)^{2}+1.$ (16)
@@ IV Optimal Sampling Temperature
IV.1 Single Gaussian Spectrum
@@
Our first implementation of cross validation will be to provide support for the optimal $\Theta$ criterion, Eq. (11). The motivation behind this criterion is to place the simulation in a temperature range safely above the regime of over-fitting, which we will be able to identify by tracking the behavior of the validation $\chi^{2}$ as $\Theta$ is lowered.
@@
For our tests, we will use synthetic $G(\tau)$ data aimed to mimic the data from a QMC simulation. Given an artificial spectrum $S(\omega)$, we use Eq. (4) to calculate the exact imaginary-time correlation function. To introduce the statistical errors necessarily present in QMC data, we generate many $G(\tau)$ bins with noise that is correlated in imaginary-time. This is done by first generating normally distributed noise for each $\tau$ point $\sigma_{0}(\tau_{i})$, and then taking the weighted averaging over all $\tau$ points with an exponentially decaying weight function:
$\sigma(\tau_{i})=\sum_{j=1}^{N_{\tau}}\sigma_{0}(\tau_{j})e^{-|\tau_{i}-\tau_{j}|}.$ (17)
@@
It was shown that the presence of covariance in the data actually improves the results of the SAC method [14], so including this in our synthetic data is critical when comparing to QMC data.
@@
We use a spectrum consisting of a single Gaussian peak, centered at $\omega=3$ and with a width $\sigma=0.5$. This spectrum can be essentially perfectly reproduced using the unconstrained SAC procedure. Here, $S(\omega)$ is represented as a weighted sum of $\delta$-functions that can freely move around in frequency-space,
$S(\omega)=\sum_{i}A_{i}\delta(\omega-\omega_{i}).$ (18)
@@
In this example, we use $N_{\omega}=5,000$ total $\delta$-functions, which is large enough to no longer detect and dependance of $S(\omega)$ on $N_{\omega}$.
@@
We generate $K=20$ independent validation data sets, each containing 1000 bins of synthetic $G(\tau)$ data. While it is certainly reasonable to generate this volume of data using modern QMC methods, this would be an inefficient use of computational resources given that the majority of the data is being used for validation. However, we use this large quantity so that the simplifications made in the large $K$ limit are more accurately realized. As we show in Sec. V.1, one can still use cross validation with far fewer $G(\tau)$ bins, even when this limit is not fully reached.
@@
When converting from $S(\omega)$ to $G(\tau)$, we set the inverse temperature $\beta=2$ and use a $\tau$ spacing of $\Delta_{\tau}=0.0625$, giving $N_{\tau}=16$ imaginary time points. The relatively low value of $\beta$ used here was chosen in light of the fact that spectra consisting of broad and smooth features, such as the simple Gaussian used here, would usually be found at higher temperatures. We generate synthetic data from this spectrum with an error level $\sigma=10^{-5}$, where the error level is defined as the magnitude of the error bar on the averaged correlation function, $\bar{G}_{k}(\tau_{i})$, for the largest $\tau$ point included, when normalized so that $G_{k}(0)=1$. For larger $\beta$, a cutoff is typically introduced at the $\tau$ value where the relative error on $G(\tau)$ is greater than 10%, which we instead use to quantify the error level.
@@
Figure 2: Deviation between the exact spectrum and those sampled at four different temperatures: the temperature where $a=0.5$ (blue), the temperature where $\chi^{2}_{\mathrm{val}}$ is minimized (red), and temperatures above/below (yellow/green) this minimum. Inset: the spectral function themselves. In this case, it is difficult to distinguish the different spectra, as they nearly overlap.
@@
The results of the cross validation procedure are shown in Fig. 1 and Fig. 2. Fig. 1 show the behavior of $\chi^{2}$ (panel (a)) and $\chi^{2}_{\mathrm{val}}$ (panel (b)) versus $\Theta$, averaged over all $K=20$ validation data sets. The temperature range corresponding to Eq. (11) with $a=0.25-1.0$ is shaded in blue. The temperature where $\chi^{2}_{\mathrm{val}}$ is at its minimum is marked by a red vertical line, with the red dashed lines denoting the standard deviation of the location of the minimum. In practice, it is useful to repeat this process, rotating which of the $K+1$ data sets is used for sampling and which are used for validation, averaging the results in the end. However, for this test we found this this was not necessary given the large number of synthetic data bins that can be easily generated.
@@
In Fig. 2, we show the difference between the the exact spectrum and the SAC spectra sampled at the temperature where $a=0.5$ (blue), at the $\chi^{2}_{\mathrm{val}}$ minimum (red), and at temperatures above (yellow) and below (green) the validation minimum, for comparison. In the inset of Fig. 2, we show the spectral functions themselves. To generate these spectra we ran the SAC procedure using all bins, recombined from the $K+1=21$ sets, thus taking full advantage of the data at hand.
@@
In this case, the minimum of $\chi^{2}_{\mathrm{val}}$ is quite sharp, and agrees very well with the temperature from Eq. (11). Consequently, the red and blue spectra in Fig. 1 are nearly identical, and, for all intents and purposes, are equally valid representations of the true spectrum, shown in black. The green and yellow spectra are included to show the relatively small $\Theta$ dependence for this simple spectrum.
IV.2 Double Guassian Spectrum
We now turn to an example where the unconstrained sampling parameterization in Eq. (18) cannot reproduce the artificial spectrum quite as closely. The spectrum we consider is composed of a pair of Gaussian peaks, one broader than the other, that intersect slightly. The area near the sharp dip, where the two peaks intersect, poses difficulties and would require a lower error level than one could reasonably achieve using real QMC data to fully resolve [14]. As such, the criterion in Eq. (11) provides an estimate for the temperature corresponding to the best possible spectrum, given these limitations.

### v2

@@ Abstract
Stochastic Analytic Continuation (SAC) of Quantum Monte Carlo (QMC) imaginary-time correlation function data is a valuable tool in connecting many-body models to experimentally measurable dynamic response functions. Recent developments of the SAC method have allowed for spectral functions with sharp features, e.g. narrow peaks and divergent edges, to be resolved with unprecedented fidelity. Often times, it is not known what exact sharp features, if any, are present a priori, and, due to the ill-posed nature of the analytic continuation problem, multiple spectral representations may be acceptable. In this work, we borrow from the machine learning and statistics literature and implement a cross validation technique to provide an unbiased method to identify the most likely spectrum amongst a set obtained with different spectral parameterizations and imposed constraints. We demonstrate the power of this method with examples using imaginary-time data generated by QMC simulations and synthetic data generated from artificial spectra. Our procedure, which can be considered a form of model selection, can be applied to a variety of numerical analytic continuation methods, beyond just SAC.
@@ I Introduction
Dynamic response functions (spectral functions) provide a key link between the quantum Monte Carlo (QMC) simulations of many-body systems and their experimental counterparts. While it is intractable to measure these spectral function directly, one can instead preform numerical analytic continuation of imaginary-time correlation functions to real frequency. However, this process is what is considered an “ill-posed” problem [1, 2, 3], meaning that there may be many viable solutions for the spectral function fitting a set of QMC-generated data. The Maximum Entropy method (MEM) [4, 5, 6, 7] has been the traditional tool used to tackle this problem and has produced many important results [8]. However, MEM has its limitations, being able only to reproduce broad spectral features, unless prior information about more complicated structure is provided to the algorithm.
@@
An alternative approach to MEM is to instead rely on the stochastic averaging of many solutions for the spectral function that fit the QMC-generated data well. The Stochastic Analytic Continuation (SAC) method [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 18, 20, 21], also referred to as the average spectrum method, achieves this by regularizing the average spectra using a fictitious temperature, $\Theta$, which mediates the competition between the minimization of the goodness-of-fit $\chi^{2}$ with respect to the QMC-generated data and the favoring of spectra with the high configurational entropy. It is now known that this procedure is equivalent to the MEM when the number of sampled degrees of freedom of the spectrum is large and when the two resulting spectra are compared at the same $\chi^{2}$ value (although the definition of the entropy must be adjusted depending on what degrees of freedom are sampled) [11, 20, 21]. Recent developments of the SAC method have also allowed for the resolution of spectra with sharp features, such as narrow quasi-particle peaks and edges with power-law singularities, by imposing appropriate constraints on the sampling space of the spectral function [16, 20].
@@
In both the plain, “unconstrained” version of SAC, and the newly developed “constrained” SAC sampling schemes, the average $\chi^{2}$ of the sampled spectra is used to optimize various parameters and gauge the ability of the spectral representation to describe the data. While these methods that rely on the average $\chi^{2}$ alone have been shown to produce reliable and consistent results, the goal of this study is to provide further predictive power for these simple schemes by taking advantage of concepts from the machine learning and statistics literature. We do this by implementing a cross validation procedure [22, 23].
@@
Cross validation is a tool used to compare how well different models describe a set of data. This is generally achieved by splitting the available data into mutually exclusive sets; a training, or sampling, set and a series of testing, or validation, sets. A model is trained on the sampling set, typically by optimizing some type of loss function. After training, the ability of the optimized model to describe the validation data sets is determined using this same loss function. This can provide information about the suitability of the model to describe the data, as well as the level of over-fitting to noise in the sampling data set.
@@
Cross validation has a natural extension to SAC. The imaginary-time correlation function data, calculated using QMC simulations, are typically stored as individual bins, which can be split up into sampling and validation sets. Here, the goodness-of-fit $\chi^{2}$ is used as the loss function to “train” and to “validate” the resulting spectral function. The concept of using cross validation within SAC was first explored by Efremkin et al. [24], whose work we draw on and significantly expand upon. Here, we preform an exhaustive study using both synthetic and real QMC-generated data.
@@
In Sec. II we describe the SAC procedure generally, and in Sec. III we explain our cross validation scheme in detail. In Sec. IV.1 and  IV.2 we implement cross validation using the unconstrained SAC method and demonstrate how it can be used to determine the optimal sampling temperature, which we compare to the simple criterion proposed previously [20]. In Sec. V.1 and  V.2, we show how cross validation can be used in practice to determine which unconstrained SAC sampling parameterization produces the most statistically likely spectrum. We extend this model selection procedure to the class of constrained SAC parameterization in Sec. VI. In Sec. VI.1 we preform a test using QMC-generated data for the $S=1/2$ antiferromagnetic (AFM) Heisenberg model, where it is known that the spectral function for the operator $O=S_{q}^{z}$ contains a power-law divergent edge. In Sec. VI.2 we preform a preliminary test on the $S=1/2$ AFM Heisenberg model with long-range interactions [25, 26], a model where the exact features of the spectral function are unknown. We summarize our results and discuss the remaining open questions and future directions in Sec. VII.
@@ II SAC Procedure
We consider the imaginary-time correlation function of an operator $O$ at inverse temperature $1/T=\beta$, evaluated at time points $\tau\in\left[0,\beta/2\right]$,
@@
Here we consider only bosonic operators, but this procedure can be extended to fermionic operators as well, with only minor adjustments to the SAC procedure. The imaginary-time dependence is defined in the Heisenberg picture, taking $\hbar=1$,
@@
The corresponding spectral function can be defined in terms of the eigenstates $|n\rangle$ and energy eigenvalues $E_{n}$ of the Hamiltonian describing the system,
@@
For the cross validation procedure, we consider $N_{B}$ bins of imaginary-time correlation function data evaluated at a set of $N_{\tau}$ points $\tau_{i}$, $\{G^{b}(\tau_{i})\}$, where $b=1,2,\ldots,N_{B}$ and $i=1,2,\ldots,N_{\tau}$. The $N_{B}$ bins are then split into $K+1$ mutually exclusive sets $\{G^{b}(\tau_{i})\}_{k}$ of equal length $N_{k}={N_{B}}/\left({K+1}\right)$. For a single cross validation run, one of the $K+1$ groups will act as the sampling data set, while the other $K$ groups will be used for cross validation. For each group, we compute the average
@@
We define the “error level” for these sets of $G(\tau)$ data as the magnitude of the error bar on the averaged correlation function, $\bar{G}_{k}(\tau_{i})$, for the largest $\tau$ point included, when normalized so that $\bar{G}_{k}(0)=G^{b}(0)=1$. For larger $\beta$, a cutoff is typically introduced at the $\tau$ value where the relative error on $\bar{G}_{k}(\tau)$ is greater than 10%, which we would instead use to quantify the error level.
The error level provides a rough gauge of the data quality, but because the statistical errors of different $\tau$ points are correlated, the full characterization of the error, and thus the calculation of $\chi^{2}$, requires the covariance matrix [7]. We use the bootstrap method to calculate the covariance matrices for each set,
$C_{k,ij}=\frac{1}{M}\sum_{m=1}^{M}(G_{k}^{m}(\tau_{i})-\bar{G}_{k}(\tau_{i}))(G_{k}^{m}(\tau_{j})-\bar{G}_{k}(\tau_{j}))$ (7)
where $G_{k}^{m}(\tau_{i})$ is the average of a bootstrap sample of $N_{k}$ randomly chosen bins among the $N_{k}$ bins in the $k$th data set and $M$ is the total number of bootstrap samples, typically much larger than $N_{B}$.
Given $S(\omega)$, the corresponding $G(\tau_{i})$ values are computed according to Eq. (4) and the $\chi^{2}$ with respect to the QMC-generated data is given by
$\chi^{2}_{k}=\sum_{i,j}^{N_{\tau}}[G(\tau_{i})-\bar{G}_{k}(\tau_{i})]\left[C_{k}^{-1}\right]_{ij}[G(\tau_{j})-\bar{G}_{k}(\tau_{j})].$ (8)
Alternatively, one can compute $\chi^{2}_{k}$ in the eigenbasis of the covariance matrix:
@@
where $\left(\sigma_{k}(\tau_{i})\right)^{2}$ are the eigenvalues of the covariance matrix and both $G(\tau_{i})$ and $\bar{G}_{k}$ have now been transformed to this eigenbasis.
In a SAC run, the spectrum $S(\omega)$ is parameterized as a large number of $\delta$-functions in a frequency continuum (in practice a very dense frequency grid). As will be described in detail later, the amplitudes and positions of the collection of $\delta$-functions can be subject to constraints to produce different spectral forms. The parameterizations (i.e. constraints) considered here are depicted in Fig. 1.
Figure 1: Depictions of the parameterizations of the spectrum $S(\omega)$: (a) Equal amplitudes and sampled frequencies in the continuum. (b) Sampled amplitudes and sampled frequencies in the continuum. (c) A “macroscopic” $\delta$-function with amplitude $A_{0}$ at $\omega_{0}$, followed by a large number $N_{\omega}$ of “continuum” $\delta$-functions at positions $\omega_{i}>\omega_{0}$ with equal amplitudes $A_{i}=(1-A_{0})/N_{\omega}$. The amplitude $A_{0}$ is fixed in a sampling run, but the frequency $\omega_{0}$ is sampled. (d) Equal amplitudes with monotonically increasing spacing $d_{i}=\omega_{i+1}-\omega_{i}$. Just as in (b), the lowest frequency $\omega_{1}$ is sampled along with all other frequencies, with the constraint $d_{i+1}>d_{i}$. (e) Varying amplitudes with monotonically increasing spacing, as in (c). In all cases, the final spectrum is the mean amplitude density of the $\delta$-functions, which is accumulated in a histogram during the sampling process.
The configurations of the $\delta$-functions are importance sampled with a weight
@@
The fictitious temperature $\Theta$ is gradually reduced, as in simulated annealing, until a minimum is reached. In the limit $\Theta\to 0$, the sampled spectrum purely minimizes $\chi^{2}$, which will not reproduce the most probable spectrum when noisy data are used. Purely Bayesian arguments imply that the temperature should be fixed to $\Theta=1$ [9, 14]. However, for $S(\omega)$ parameterized by a large number of sampled degrees of freedom, entropy will ruin the goodness-of-fit at this fixed value of $\Theta$ [20]. Instead, it has been proposed to use a temperature determined by a simple criterion to give the optimal balance of entropy and goodness-of-fit, corresponding to the value where
$\langle\chi^{2}(\Theta)\rangle=\chi^{2}_{\rm min}+a\sigma_{\chi^{2}},$ (11)
where $\sigma_{\chi^{2}}=\sqrt{2\chi^{2}_{\rm min}}$ is used as a proxy for the standard deviation of the $\chi^{2}$ distribution that the value of $\langle\chi^{2}\rangle$ follows and $a$ is an order-one number (typically $a=1/2$ is used). This criterion, motivated by the properties of the $\chi^{2}$ distribution, has been shown to produce reliable results, reproducing spectra generated from synthetic QMC $G(\tau)$ data with high fidelity [20].
@@ III Cross Validation Procedure
which can be split into three terms:
@@
$X_{3}=\frac{2}{K}\sum_{k=1}^{K}\sum_{i=1}^{N_{\tau}}\frac{\left(\bar{G}_{k}(\tau_{i})-G_{\rm ex}(\tau_{i})\right)\left({G}_{\rm ex}(\tau_{i})-{G}(\tau_{i})\right)}{\sigma^{2}_{k}(\tau_{i})}.$ (15c)
Here $X_{1}$ contains purely information about how well the estimated $G(\tau)$, calculated from the sampled $S(\omega)$, describes the exact correlation function. This will vary as we change $\Theta$, or any other parameter used in the SAC program, and will be minimized by the most statistically likely spectral function, given input data $\bar{G}_{0}(\tau)$ and the corresponding covariance matrix $C_{0}$.
$X_{2}$ is the the $\chi^{2}$ value for the validation data with respect to the underlying data $G_{\rm ex}$, averaged over each of the $K$ validation data sets. The value of $X_{2}$ will follow the $\chi^{2}$ distribution with $N_{\rm dof}=N_{\tau}$ and will have a mean of $E\left[X_{2}\right]=N_{\rm dof}=N_{\tau}$ and a variance of $V\left[X_{2}\right]=2N_{\rm dof}=2N_{\tau}$. This term is independent of any SAC parameter and will simply contribute a background value of $N_{\tau}$, in the limit of large $K$.
$X_{3}$ contains both information about the accuracy of $G(\tau)$ with respect to ${G}_{\rm ex}(\tau)$ and the deviation of the $\bar{G}_{k}(\tau)$ with respect to ${G}_{\rm ex}(\tau)$. However, these two factors are uncorrelated, and since we expect that the fluctuations of $\bar{G}_{k}(\tau)$ will follow a Gaussian distribution with a mean of zero, this term will also have a mean of zero.
@@
$\frac{\chi^{2}_{\mathrm{val}}}{N_{\tau}}\coloneqq\frac{1}{N_{\tau}K}\sum_{k=1}^{K}\sum_{i=1}^{N_{\tau}}\left(\frac{G(\tau_{i})-{G}_{\rm ex}(\tau_{i})}{\sigma_{k}(\tau_{i})}\right)^{2}+1.$ (16)
@@ IV Optimal Sampling Temperature
IV.1 Single Gaussian Peak
@@
Our first implementation of cross validation will be to provide support for the optimal $\Theta$ criterion, Eq. (11). The motivation behind this criterion is to achieve a statistically good data fit while placing the simulation in a temperature range safely above the regime of over-fitting, which we will be able to identify by tracking the behavior of the validation $\chi^{2}$ as $\Theta$ is lowered.
@@
For our tests, we will use synthetic $G(\tau)$ data aimed to mimic the data generated in a QMC simulation. Given an artificial spectrum $S(\omega)$, we use Eq. (4) to calculate the exact imaginary-time correlation function. To introduce the statistical errors necessarily present in QMC-generated data, we generate many $G(\tau)$ bins with noise that is correlated in imaginary-time. This is done by first generating normally distributed noise for each $\tau$ point $\varepsilon_{0}(\tau_{i})$, and then taking the weighted averaging over all $\tau$ points with an exponentially decaying weight function:
$\varepsilon(\tau_{i})=\sum_{j=1}^{N_{\tau}}\varepsilon_{0}(\tau_{j})e^{-|\tau_{i}-\tau_{j}|}.$ (17)
@@
It was shown that the presence of covariance in the data actually improves the results of the SAC method [20], so including this in our synthetic data is critical when comparing to QMC-generated data.
@@
In our first test, we consider a spectrum consisting of a single Gaussian peak, centered at $\omega=3$. This spectrum can be essentially perfectly reproduced using the unconstrained SAC procedure, Fig. 1(a) and (b). Here, $S(\omega)$ is represented as a sum of $N_{\omega}$ $\delta$-functions that can freely move around in frequency-space,
$S(\omega)=\sum_{i}A_{i}\delta(\omega-\omega_{i}),$ (18)
@@
in some cases with $A_{i}$ variable and sampled [11], as depicted in Fig. 1(b), but in this case we use fixed amplitudes $A_{i}=1/N_{\omega}$, Fig. 1(a). In this example, we use $N_{\omega}=5,000$ total $\delta$-functions, which is large enough to no longer detect and dependance of $S(\omega)$ on $N_{\omega}$.
@@
We generated $K=20$ independent validation data sets, each containing 1,000 bins of synthetic $G(\tau)$ data with an error level $\sigma=10^{-5}$. While it is certainly reasonable to generate this volume of data using modern QMC methods, this would be an inefficient use of computational resources given that the majority of the data is being used for validation. However, we used this large quantity so that the simplifications made in the large $K$ limit are more accurately realized. As we show in Sec. VI, one can still use cross validation with far fewer $G(\tau)$ bins, even when this limit is not fully reached.
@@
When converting from $S(\omega)$ to $G(\tau)$, we set the inverse temperature $\beta=2$ and use a $\tau$ spacing of $\Delta_{\tau}=0.0625$, giving $N_{\tau}=16$ imaginary time points. The relatively low value of $\beta$ used here was chosen in light of the fact that spectra consisting of broad and smooth features, such as the simple Gaussian used here, would usually be found at higher temperatures.
@@
Figure 3: Deviation between the exact spectrum and those sampled at four different temperatures $\Theta$: that where $a=0.5$ (blue), where $\chi^{2}_{\mathrm{val}}$ is minimized (red), and $\Theta$ values above/below (yellow/green) this minimum. Inset: the spectral function themselves. In this case, it is difficult to distinguish the different spectra, as they nearly overlap.
@@
The results of the cross validation procedure are presented in Fig. 2 and Fig. 3. Fig. 2 show the behavior of the sampling $\chi^{2}$ (panel (a)) and the validation $\chi^{2}$ (panel (b)) versus $\Theta$, averaged over all $K=20$ validation data sets and normalized by the number of $\tau$ points. The temperature range corresponding to Eq. (11) with $a=0.25-1.0$ is shaded in blue. The temperature where the validation $\chi^{2}$ is at its minimum is marked by a red vertical line, with the red dashed lines denoting the standard deviation of the location of the minimum. In practice, it is useful to repeat this process, rotating which of the $K+1$ data sets is used for sampling and which are used for validation, and averaging the results in the end. However, for this test we found this this was not necessary given the large number of synthetic data bins that can be easily generated.
@@
In Fig. 3, we show the difference between the the exact spectrum and the SAC spectra sampled at the temperature where $a=0.5$ (blue), at the validation $\chi^{2}$ minimum (red), and at temperatures above (yellow) and below (green) the validation minimum, for comparison. In the inset of Fig. 3, we show the spectral functions themselves. To generate these spectra we ran the SAC procedure using all bins, recombined from the $K+1=21$ sets, thus taking full advantage of the data at hand.
@@
In this case, the minimum of the validation $\chi^{2}$ is quite sharp and agrees very well with the temperature from Eq. (11). Consequently, the red and blue spectra in Fig. 2 are nearly identical, and, for all intents and purposes, are equally valid representations of the true spectrum, shown in black. The green and yellow spectra are included to show the relatively small $\Theta$ dependence for this simple spectrum.